# Do global forecasting models require frequent retraining?

### v1

@@ Acknowledgements
Figure 9: VN1 Friedman-Nemenyi test results based on MQL.
The costs tables 9 and 10 show the estimated cost in real values of each scenario for daily and weekly data respectively.
Method 7 14 21 30 60 90 120 150 180 364 LR 390,732\$ 208,499\$ 150,791\$ 112,547\$ 69,132\$ 52,131\$ 48,474\$ 46,678\$ 38,937\$ 33,151\$ XGBoost 529,679\$ 281,358\$ 200,778\$ 151,635\$ 88,200\$ 69,495\$ 60,082\$ 60,155\$ 48,189\$ 39,861\$ LGBM 1,526,424\$ 818,569\$ 593,676\$ 455,075\$ 286,262\$ 228,698\$ 199,337\$ 204,972\$ 170,802\$ 146,252\$ CatBoost 358,123\$ 189,714\$ 134,258\$ 99,699\$ 58,167\$ 44,164\$ 37,206\$ 38,266\$ 30,622\$ 24,362\$ MLP 604,120\$ 482,505\$ 441,981\$ 419,823\$ 392,046\$ 383,287\$ 380,697\$ 376,968\$ 368,023\$ 365,525\$ TCN 1,146,256\$ 960,626\$ 890,924\$ 862,432\$ 815,006\$ 801,612\$ 791,801\$ 791,348\$ 784,201\$ 776,181\$ NBEATSx 729,263\$ 561,105\$ 514,166\$ 466,683\$ 425,593\$ 411,232\$ 400,953\$ 401,352\$ 393,673\$ 387,058\$ NHITS 754,783\$ 573,233\$ 510,655\$ 470,077\$ 425,684\$ 409,844\$ 402,731\$ 402,394\$ 393,071\$ 385,470\$ Average 754,922\$ 509,451\$ 429,654\$ 379,746\$ 320,011\$ 300,058\$ 290,160\$ 290,267\$ 278,440\$ 269,733\$
Table 9: M5 estimated costs for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52 LR 15,234\$ 7,839\$ 5,731\$ 4,292\$ 3,205\$ 2,508\$ 2,140\$ 2,190\$ 1,463\$ 1,099\$ RF 1,605,768\$ 799,597\$ 562,896\$ 398,954\$ 281,246\$ 199,214\$ 158,959\$ 165,975\$ 81,529\$ 38,865\$ XGBoost 34,254\$ 17,687\$ 12,796\$ 9,413\$ 6,944\$ 5,262\$ 4,473\$ 4,534\$ 2,863\$ 2,005\$ LGBM 505,304\$ 260,721\$ 186,904\$ 138,460\$ 101,775\$ 76,738\$ 65,013\$ 66,335\$ 40,721\$ 27,701\$ CatBoost 52,021\$ 26,594\$ 18,515\$ 13,366\$ 9,611\$ 7,022\$ 5,798\$ 5,977\$ 3,311\$ 2,023\$ MLP 62,102\$ 34,489\$ 26,431\$ 21,149\$ 17,121\$ 14,391\$ 13,027\$ 13,015\$ 10,255\$ 8,891\$ LSTM 82,927\$ 49,019\$ 38,851\$ 31,962\$ 27,022\$ 23,465\$ 21,829\$ 21,932\$ 18,415\$ 16,652\$ TCN 72,763\$ 41,291\$ 31,910\$ 25,464\$ 20,791\$ 17,579\$ 15,983\$ 16,365\$ 12,936\$ 11,284\$ NBEATSx 80,337\$ 44,837\$ 34,283\$ 27,179\$ 22,057\$ 18,387\$ 16,686\$ 16,702\$ 13,085\$ 11,277\$ NHITS 80,776\$ 45,458\$ 34,688\$ 27,585\$ 22,336\$ 18,607\$ 16,888\$ 16,855\$ 13,405\$ 11,601\$ Average 259,149\$ 132,753\$ 95,301\$ 69,782\$ 51,211\$ 38,317\$ 32,080\$ 32,988\$ 19,798\$ 13,140\$
Table 10: VN1 estimated costs for each method and retrain scenario combination.

### v2

@@ 5 Conclusions
Figure 11: VN1 Friedman-Nemenyi test results based on SMQL.
Method 7 14 21 30 60 90 120 150 180 364 LR 390,732 208,499 150,791 112,547 69,132 52,131 48,474 46,678 38,937 33,151 XGBoost 529,679 281,358 200,778 151,635 88,200 69,495 60,082 60,155 48,189 39,861 LGBM 1,526,424 818,569 593,676 455,075 286,262 228,698 199,337 204,972 170,802 146,252 CatBoost 358,123 189,714 134,258 99,699 58,167 44,164 37,206 38,266 30,622 24,362 MLP 604,120 482,505 441,981 419,823 392,046 383,287 380,697 376,968 368,023 365,525 TCN 1,146,256 960,626 890,924 862,432 815,006 801,612 791,801 791,348 784,201 776,181 NBEATSx 729,263 561,105 514,166 466,683 425,593 411,232 400,953 401,352 393,673 387,058 NHITS 754,783 573,233 510,655 470,077 425,684 409,844 402,731 402,394 393,071 385,470 Average 754,922 509,451 429,654 379,746 320,011 300,058 290,160 290,267 278,440 269,733
Table 9: M5 estimated costs (in \$) for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52 LR 15,234 7,839 5,731 4,292 3,205 2,508 2,140 2,190 1,463 1,099 RF 1,605,768 799,597 562,896 398,954 281,246 199,214 158,959 165,975 81,529 38,865 XGBoost 34,254 17,687 12,796 9,413 6,944 5,262 4,473 4,534 2,863 2,005 LGBM 505,304 260,721 186,904 138,460 101,775 76,738 65,013 66,335 40,721 27,701 CatBoost 52,021 26,594 18,515 13,366 9,611 7,022 5,798 5,977 3,311 2,023 MLP 62,102 34,489 26,431 21,149 17,121 14,391 13,027 13,015 10,255 8,891 LSTM 82,927 49,019 38,851 31,962 27,022 23,465 21,829 21,932 18,415 16,652 TCN 72,763 41,291 31,910 25,464 20,791 17,579 15,983 16,365 12,936 11,284 NBEATSx 80,337 44,837 34,283 27,179 22,057 18,387 16,686 16,702 13,085 11,277 NHITS 80,776 45,458 34,688 27,585 22,336 18,607 16,888 16,855 13,405 11,601 Average 259,149 132,753 95,301 69,782 51,211 38,317 32,080 32,988 19,798 13,140
Table 10: VN1 estimated costs (in \$) for each method and retrain scenario combination.
Figure 12: M5 optimal retraining frequency results for each model based on RMSSE.
Figure 13: M5 optimal retraining frequency results for each model based on SMQL.
Figure 14: VN1 optimal retraining frequency results for each model based on RMSSE.
Figure 15: VN1 optimal retraining frequency results for each model based on SMQL.